# Inference for the stochastic FitzHugh-Nagumo model from real action potential data via approximate Bayesian computation

### v1

@@ 4.2 Inference under different summary statistics
In Figure 5, we report the 90% CIs of the marginal posterior densities obtained via Algorithm 1 using the canonical (red lines) or the structure-based (black lines) summaries (the latter corresponding to the SBP SMC-ABC method). These CIs are plotted as a function of the computational budget, ranging up to $500000$ (top panels), $1500000$ (middle panels) or $6\cdot 10^{6}$ (bottom panels). For $\beta$ and $\sigma$, the CIs obtained when using the canonical summaries are wider than those derived from the structure-based summaries up to circa $2$-$3$ million model simulations, when they have approximately the same width. Moreover, using structure-based summaries yields CIs approaching the desired posterior region much faster, centered around the true parameter values for as few as $200$-$300$ thousand simulations, and remaining almost unchanged when increasing the computational budget. This is not the case when using the canonical summaries. On the one hand, they lead to posteriors with CIs for $\epsilon$ and $\sigma$ continuing shrinking when increasing the number of simulations (see the bottom panels). On the other hand, they yield posterior medians much larger than the true parameter values under small computational budgets (see the top panels), and below the true values under larger computational budgets (see in particular the bottom left panel for $\epsilon$).
@@
Figure 5: 90% CIs of the marginal posterior densities of the parameter vector $\theta$ (2) of the stochastic FHN model (1) obtained from Algorithm 1 with either canonical (cf. Remark 2, red lines) or structure-based (cf. (7), black lines) summaries as a function of the computational budget. The green horizontal lines and the dashed lines indicate the true parameter values and the corresponding posterior medians, respectively.
Figure 6: Marginal posterior densities of the parameter vector $\theta$ (2) of the stochastic FHN model (1) obtained from the SMC-ABC Algorithm 1 with either canonical (cf. Remark 2, red lines) or structure-based (cf. (7), black lines) summaries when using approximately $3\cdot 10^{5}$ (dashed lines) and $11$-$13$ million (solid lines) model simulations, respectively. The green vertical lines and the blue horizontal lines indicate the true parameter values and the prior densities, respectively.
@@
Similar observations can be made from Figure 6, where we plot the marginal posterior densities for a computational budget of approximately $3\cdot 10^{5}$ (dashed lines) and 12 million (solid lines) model simulations. While the structure-based summaries (black lines) yield satisfactory results even under a small computational budget, the canonical summaries (red lines) fail to do so. The results for $\epsilon$ are particularly problematic, suggesting that the canonical summaries may be driving the posterior mass to a narrow and “wrong” area when increasing the computational budget, as also observed from the bottom left panel in Figure 5.
@@ 5.1 Description of the data
The investigated data have been obtained via a neural recording from the 5th lumbar dorsal rootlet of an adult female Sprague Dawley rat, and are available at https://data.mendeley.com/datasets/ybhwtngzmm/1. A detailed description of the data and the corresponding experimental design are provided in [40]. In brief, these data consist of twenty recordings of length $T_{\textrm{obs}}=250$ ms with a sampling rate of $50$ kHz (and thus $\Delta_{\textrm{obs}}=0.02$), ten obtained with the animal at rest (see the datasets in the folder “Resting”), and ten during stimulation (of the dermatome, activating the corresponding axons in the dorsal root, see the datasets in the folder “Cutaneous Stimulation”). Each recording consists of $5$ channels of voltage data.
@@ 5.2 Parameter inference
Figure 7: Marginal posterior densities of the parameter vector $\theta$ (2) of the stochastic FHN model (1) fitted on the real action potential datasets 1553 (animal at rest, black solid lines), 1554 (animal stimulated, red solid lines), 1608 (animal at rest, black dashed lines), and 1609 (animal stimulated, red dashed lines), using the SBP SMC-ABC method. The horizontal blue lines indicate the uniform prior densities (11).
@@
ABC setup
The stochastic FHN model (1) is fitted to these four datasets using the SBP SMC-ABC method, with a computational budget of $\textrm{Nsim}_{\textrm{max}}=2\cdot 10^{6}$. All synthetic datasets are generated with step-size $\Delta=\Delta_{\textrm{obs}}=0.02$, and time horizon $T=T_{\textrm{obs}}=250$. Note that, the real datasets are centered close to zero, with a mean value of approximately $0.003$ (the empirical means of the datasets 1553, 1608, 1554 and 1609 are $0.00277$, $0.00265$, $0.00271$ and $0.00263$, respectively). Therefore, in the following, all investigated datasets (real and synthetic) are centered.
Estimation results
All marginal posterior densities are unimodal, notably different from the priors. Interestingly, the posteriors for the two “resting” (respectively “stimulating”) scenarios are similar, while they differ across these two cases, with the ABC marginal posterior densities of all parameters shifted towards larger values in the “stimulating” scenarios compared to the “resting” scenarios (see also Table 2). Moreover, their uncertainty also increases under the “stimulating” settings, as confirmed by the larger posterior standard deviations, and the wider $90$% CIs reported in Table 2.
Fitted FHN model
Fitting real data, we do not have ground truth parameters to compare our ABC results with. However, we check here whether the obtained results fit the model, which would advocate in favour of both the model and the inferential procedure. Figure 8 reports the estimated structure-based summary statistics (7) from the four investigated real action potential recordings (black solid lines) in comparison with the summaries estimated from synthetic datasets simulated under fifty parameter values sampled from the ABC posterior according to their weights (gray areas), and under the ABC posterior means reported in Table 2 (red dashed lines).
Figure 9: Exemplary 30 ms data snippet from recording 1553 (top panel) compared to a simulated dataset obtained under the corresponding ABC posterior means (middle panel). Exemplary real action potential (bottom left panel) taken from recording 1553 compared to a simulated action potential (bottom right panel) taken from the synthetic dataset visualised in the middle panel.
This nice fit is also observed in Figure 9. There, we compare both an exemplary $30$ ms snippet from the real action potential recording 1553 (top panel) with a $30$ ms simulated dataset generated under the corresponding posterior means reported in Table 2 (middle panel), and an action potential taken from the real data recording 1553 (bottom left panel) with another taken from the reported simulated dataset (bottom right panel). The synthetic action potential closely resembles the shape of the real action potential, including the depolarization (rapid increase of voltage), repolarization (rapid decrease of voltage), and refractory (time during which the cell cannot repeat an action potential) periods.
@@ 6 Conclusion
We propose a SMC variant of the SBP ABC method initially presented in [6] for a different class of SDEs, with the goal of estimating the parameters of the stochastic FHN model. This is a well known SDE used to describe the spiking activity of a single neuron at the intracellular level. The proposed SBP SMC-ABC method relies on summary statistics based on the ergodicity property of the model [31], and on a recently proposed structure-preserving numerical splitting scheme for reliable and efficient synthetic data generation [5].
@@
While existing estimation methods often require complete observations, a non-degenerate noise structure, globally Lipschitz coefficients or a very complicated architecture, our method overcomes the associated difficulties, is easy-to-grasp, and straight-forward to implement and to parallelise. Moreover, in contrast to other inference procedures (which, e.g. require to fix the time scale separation parameter $\epsilon$, see [9]), the SMC SBP-ABC algorithm succeeds in simultaneously estimating all four model parameters of the stochastic FHN model, even for fewer observations and relatively small observation time horizons.
@@
While there exist very few real data studies using the FHN model, the presented analysis of real action potential is the first for the considered stochastic version of the model. Here, we investigate real data recorded under different experimental conditions (resting and cutaneous stimulation) in an adult female rat [40]. The inferred posterior distributions are in agreement with the different experimental settings, and lead to satisfactory goodness-of-fit plots, suggesting the goodness of the considered model and the proposed estimation approach.
@@
Besides successfully deriving marginal posterior densities from both synthetic and real data, we also discuss the influence of different key ingredients of ABC. For example, the proposed structure-based summaries avoid possible biases which may be instead introduced by the canonical ones, while other choices of percentiles (e.g. the 25th) or distances (e.g. the Wasserstein) have a limited impact on the results. In terms of proposal samplers, which one to choose depends on the available computational budget. Indeed, olcm (and other guided sequential ABC schemes proposed in [46]) leads to more accurate posterior inference than standard (the perturbation kernel considered here) for low/moderate computational budgets, with the differences becoming negligible as the computational budget increases. On the other hand though, olcm and the guided sequential schemes are characterised by more involved proposal means and covariances, introducing a computational overhead with respect to standard.
@@
Finally, we note that the proposed method can be applied to any other ergodic SDE for which a computationally efficient and reliable numerical method for synthetic data generation is available. The exploration of more complex stochastic neuronal models is certainly of interest, and will be considered in future works.
Appendix A Impact of the proposal kernel

### v2

@@ 4.2 Inference under different summary statistics
In Figure 6, we report the 90% CIs of the marginal posterior densities obtained by applying Algorithm 1 to the observed data visualized in Figure 2 (observation setting $T_{\textrm{obs}}=200$, $\Delta_{\textrm{obs}}=0.02$), using the canonical (red lines) or the structure-based (black lines) summaries (the latter corresponding to the SBP SMC-ABC method). These CIs are plotted as a function of the computational budget, ranging up to $5\cdot 10^{5}$ (top panels), $1.5\cdot 10^{6}$ (middle panels) or $6\cdot 10^{6}$ (bottom panels). For $\beta$ and $\sigma$, the CIs obtained when using the canonical summaries are wider than those derived from the structure-based summaries up to circa $2$-$3$ million model simulations, when they have approximately the same width. Moreover, using structure-based summaries yields CIs approaching the desired posterior region much faster, centered around the true parameter values for as few as $200$-$300$ thousand simulations, and remaining almost unchanged when increasing the computational budget. This is not the case when using the canonical summaries. On the one hand, they lead to posteriors with CIs for $\epsilon$ and $\sigma$ continuing shrinking when increasing the number of simulations (see the bottom panels). On the other hand, they yield posterior medians much larger than the true parameter values under small computational budgets (see the top panels), and below the true values under larger computational budgets (see in particular the bottom left panel for $\epsilon$).
@@
Figure 6: 90% CIs of the marginal posterior densities of the parameter vector $\theta$ (2) of the stochastic FHN model (1) obtained from Algorithm 1 with either structure-based (cf. (8), black lines) or canonical (cf. Remark 2, red lines) summaries as a function of the computational budget. The green horizontal lines and the dashed lines indicate the true parameter values and the corresponding posterior medians, respectively.
@@
Similar observations can be made from Figure 7, where we plot the marginal posterior densities for a computational budget of approximately $3\cdot 10^{5}$ (dashed lines) and 12 million (solid lines) model simulations. While the structure-based summaries (black lines) yield satisfactory results even under a small computational budget, the canonical summaries (red lines) fail to do so. The results for $\epsilon$ are particularly problematic, suggesting that the canonical summaries may be driving the posterior mass to a narrow and “wrong” area when increasing the computational budget, as also observed from the bottom left panel in Figure 6 and the ABC posterior means and standard deviations reported in Table 2.
Table 2: ABC posterior means and standard deviations of $\theta$ (2), obtained from the SMC-ABC Algorithm 1 with either structure-based (cf. (8)) or canonical (cf. Remark 2) summaries when using approximately $11-13$ million model simulations.
Summaries ABC posterior means ABC posterior stand. dev.s
Structure-based $(0.100,1.509,0.797,0.297)$ $(0.007,0.069,0.050,0.017)$
Canonical $(0.097,1.482,0.792,0.292)$ $(0.002,0.051,0.042,0.019)$
Figure 7: Marginal posterior densities of the parameter vector $\theta$ (2) of the stochastic FHN model (1) obtained from the SMC-ABC Algorithm 1 with either structure-based (cf. (8), black lines) or canonical (cf. Remark 2, red lines) summaries when using approximately $3\cdot 10^{5}$ (dashed lines) and $11$-$13$ million (solid lines) model simulations, respectively. The green vertical lines and the blue horizontal lines indicate the true parameter values and the prior densities, respectively.
In Appendix A, we also investigate the computational cost of the summaries, both in terms of evaluation time per dataset (in the order of milliseconds) and runtimes of the corresponding algorithms as functions of the computational budget. Calculating canonical summaries for a given dataset is slightly faster than computing the structure-based ones, but the overall impact on the runtime is negligible. Considering that the algorithm based on structure-based summaries requires fewer model simulations to achieve a desired accuracy, these summaries are then preferable to the canonical ones also in terms of overall algorithm runtimes.
Remark 4.
Interestingly, considering the alternative structure-based summaries $s(\tilde{y}_{\theta})=(\hat{f}_{\tilde{y}_{\theta}},\hat{R}_{\tilde{y}_{\theta}})$ obtained by replacing the spectral density with the underlying auto-correlation function (here estimated for time-lags $1,\ldots,25$) does not yield inferential results as good as those obtained with the proposed structure-based summaries (8). While the resulting ABC posterior distributions also cover the true parameter values well, their standard deviations are slightly larger and their means slightly more off the true values compared to the results obtained via (8) (figures not shown). Considering $s(\tilde{y}_{\theta})=(\hat{f}_{\tilde{y}_{\theta}},\hat{S}_{\tilde{y}_{\theta}},\hat{R}_{\tilde{y}_{\theta}})$ (i.e., adding the auto-correlation function to (8)) does not yield a notable improvement. This may be explained by the fact that the spectral density already uses the information contained in the auto-correlation function.
@@ 5.1 Description of the data
The investigated data have been obtained via a neural recording from the 5th lumbar dorsal rootlet of an adult female Sprague Dawley rat, and are available at https://data.mendeley.com/datasets/ybhwtngzmm/1. A detailed description of the data and the corresponding experimental design are provided in [42]. In brief, these data consist of twenty recordings of length $T_{\textrm{obs}}=250$ ms with a sampling rate of $50$ kHz (and thus $\Delta_{\textrm{obs}}=0.02$), ten obtained with the animal at rest (see the datasets in the folder “Resting”) and ten during stimulation (of the dermatome, activating the corresponding axons in the dorsal root, see the datasets in the folder “Cutaneous Stimulation”). Each recording consists of $5$ channels of voltage data.
@@ 5.2 Parameter inference
ABC setup
The stochastic FHN model (1) is fitted to these four datasets using the SBP SMC-ABC method, with a computational budget of $\textrm{Nsim}_{\textrm{max}}=2\cdot 10^{6}$. All synthetic datasets are generated with step-size $\Delta=\Delta_{\textrm{obs}}=0.02$ and time horizon $T=T_{\textrm{obs}}=250$. Note that, the real datasets are centered close to zero, with a mean value of approximately $0.003$ (the empirical means of the datasets 1553, 1608, 1554 and 1609 are $0.00277$, $0.00265$, $0.00271$ and $0.00263$, respectively). Therefore, in the following, all investigated datasets (real and synthetic) are centered.
Estimation results
All marginal posterior densities are unimodal, notably different from the priors. Interestingly, the posteriors for the two “resting” (resp. “stimulating”) scenarios are similar, while they differ across these two cases, with the ABC marginal posterior densities of all parameters shifted towards larger values in the “stimulating” scenarios compared to the “resting” scenarios (see also Table 3). Moreover, their uncertainty also increases under the “stimulating” settings, as confirmed by the larger posterior standard deviations and the wider $90$% CIs reported in Table 3. A deeper analysis of the derived posterior distributions is provided in Appendix D.
Figure 8: Marginal posterior densities of the parameter vector $\theta$ (2) of the stochastic FHN model (1) fitted on the real action potential datasets 1553 (animal at rest, black solid lines), 1554 (animal stimulated, red solid lines), 1608 (animal at rest, black dashed lines) and 1609 (animal stimulated, red dashed lines), using the SBP SMC-ABC method. The horizontal blue lines indicate the uniform prior densities (12).
@@
Fitted FHN model
Fitting real data, we do not have ground truth parameters to compare our ABC results with. However, we check here whether the obtained results fit the model, which would advocate in favour of both the model and the inferential procedure. Figure 9 reports the estimated structure-based summary statistics (8) from the four investigated real action potential recordings (black solid lines) in comparison with the summaries estimated from synthetic datasets simulated under fifty parameter values sampled from the ABC posterior according to their weights (gray areas) and under the ABC posterior means reported in Table 3 (red dashed lines).
@@
Figure 10: Exemplary 30 ms data snippet from recording 1553 (top panel) compared to a simulated dataset obtained under the corresponding ABC posterior means (middle panel, blue solid line) and to the posterior predictive band (middle panel, shaded gray area) based on $500$ synthetic datasets simulated using the same seed for pseudo-random number generation and different parameter values re-sampled from the ABC posterior distribution. Exemplary real action potential (bottom left panel) taken from recording 1553 compared to a simulated action potential (bottom right panel) taken from the blue synthetic dataset visualised in the middle panel.
This nice fit is also observed in Figure 10. There, we compare an exemplary $30$ ms snippet from the real action potential recording 1553 (top panel) with a $30$ ms simulated dataset generated under the corresponding posterior means reported in Table 3 (middle panel, blue solid line) and a $30$ ms posterior predictive band (middle panel, shaded grey area) of synthetic datasets simulated under 500 kept ABC posterior samples (re-sampled using the corresponding weights). Moreover, we compare an action potential taken from the real data recording 1553 (bottom left panel) with another taken from the reported simulated dataset based on the ABC posterior means (bottom right panel). The synthetic action potential closely resembles the shape of the real action potential, including the depolarization (rapid increase of voltage), repolarization (rapid decrease of voltage) and refractory (time during which the cell cannot repeat an action potential) periods. Not all kept posterior samples yield paths with action potentials arising at the same times as those obtained when simulating from the ABC posterior means (blue line in the middle panel of Figure 10), which explains the slightly broader gray band around the spiking times. However, those paths contain at least one of the two exemplified action potentials and/or show spikes at other epochs, so the whole ABC posterior distribution reproduces the desired neural activity, with posterior predictive densities and spectral densities nicely matching those derived from the real dataset (cf. Figure 9).
@@ 6 Conclusion
We propose a SMC variant of the SBP ABC method initially presented in [6] for a different class of SDEs, with the goal of estimating the parameters of the stochastic FHN model. This is a well known SDE used to describe the spiking activity of a single neuron at the intracellular level. The proposed SBP SMC-ABC method relies on summary statistics based on the ergodicity property of the model [33] and on a recently proposed structure-preserving numerical splitting scheme for reliable and efficient synthetic data generation [5].
@@
While existing estimation methods often require complete observations, a non-degenerate noise structure, globally Lipschitz coefficients or a very complicated architecture, our method overcomes the associated difficulties, is easy-to-grasp and straight-forward to implement and to parallelise. Moreover, in contrast to other inference procedures (which, e.g. require to fix the time scale separation parameter $\epsilon$, see [10]), the SMC SBP-ABC algorithm succeeds in simultaneously estimating all four model parameters of the stochastic FHN model, even for fewer observations and relatively small observation time horizons.
@@
While there exist very few real data studies using the FHN model, the presented analysis of real action potential is the first for the considered stochastic version of the model. Here, we investigate real data recorded under different experimental conditions (resting and cutaneous stimulation) in an adult female rat [42]. The inferred posterior distributions are in agreement with the different experimental settings and lead to satisfactory goodness-of-fit plots, suggesting the goodness of the considered model and the proposed estimation approach.
@@
Besides successfully deriving marginal posterior densities from both synthetic and real data, we also discuss the influence of different key ingredients of ABC. For example, the proposed structure-based summaries avoid possible biases which may be instead introduced by the canonical ones, while other choices of percentiles (e.g. the 25th), distances (e.g. the Wasserstein) or priors (e.g. log-normal and exponential) have a limited impact on the results. In terms of proposal samplers, which one to choose depends on the available computational budget. Indeed, olcm (and other guided sequential ABC schemes proposed in [49]) leads to more accurate posterior inference than standard (the perturbation kernel considered here) for low/moderate computational budgets, with the differences becoming negligible as the computational budget increases. On the other hand though, olcm and the guided sequential schemes are characterised by more involved proposal means and covariances, introducing a computational overhead with respect to standard.
@@
Finally, we note that the proposed method can be applied to any other ergodic SDE for which a computationally efficient and reliable numerical method for synthetic data generation is available. The exploration of more complex stochastic neuronal models via the proposed ABC methodology (or suitable adaptations thereof) is certainly of interest and will be considered in future works. Moreover, motivated by [14], investigations using path signatures as summary statistics within ABC-inference for SDE models are currently undergoing, and may lead to successful inference also for non-ergodic models and/or multivariate observed components.
Appendix A Computational cost of different summary statistics
In Section 4.2, we discussed the impact of canonical and structure-based summary statistics on the quality of the inference, observing how the latter requires a lower computational budget (in terms of model simulations) than the former to achieve a desired accuracy level. In this section, we investigate the computational cost of calculating these summaries and the impact it has in terms of runtimes of the algorithms. We used the R-package microbenchmark to measure the time it takes to compute the structure-based and canonical summary statistics for all considered datasets, repeating such operation sequentially 5,000 times per dataset. Note that the actual evaluation time, over several cores, will be lower. The median evaluation times, in microseconds $[\mu s]$, were obtained on a MacbookPro 2020 (Processor 1.4GHz QUad-Core Intel Core i5, Memory: 8 GB 2133 MhZ) and are reported in Table 4.
Interestingly, evaluating the canonical summaries is 1.61.75×1.6-1.75$\times$1.6 - 1.75 × faster than the structure-based ones for datasets with $n=625,1250,2500$ observations, but only 1.3×1.3$\times$1.3 × faster for datasets with $n=5000,10000$ observations, when comparing the median evaluation times. In particular, computing the canonical summaries for the largest dataset takes a median time of $4.428\ \mu\textrm{s}$ compared to $5.775\ \mu\textrm{s}$ for the structure-based.
The impact on the runtimes of the algorithm is negligible though, with the SMC-ABC algorithm with structure-based summaries being actually slightly faster than that with canonical summaries. This can be observed in Figure 11 (left panel), where we report the runtimes (in minutes) of the respective algorithms as functions of the computational budget for ten runs each (launched on a machine with $48$ cores at the University of Warwick). The slightly lower runtimes obtained under the structure-based summaries may be explained by the fact that fewer samples $\theta^{*}$ were immediately rejected due to being outside the prior domain (cf. Line 19 of Algorithm 1), as observed in the right panel of Figure 11. This is a further advantage of the structure-based summaries.
$T_{\textrm{obs}}$ $\Delta_{\textrm{obs}}$ n Summary statistics Median evaluation time $[\mu s]$
200 0.02 $10^{4}$ struct.-based 5775.1880
canonical 4428.1330
0.04 5000 struct.-based 3332.5880
canonical 2487.1360
0.08 2500 struct.-based 2074.2020
canonical 1297.2495
100 0.02 5000 struct.-based 3316.5125
canonical 2475.3875
0.04 2500 struct.-based 2067.1440
canonical 1293.1045
0.08 1250 struct.-based 1550.1925
canonical 965.1840
50 0.02 2500 struct.-based 2059.7680
canonical 1289.8485
0.04 2500 struct.-based 1546.2915
canonical 968.1145
0.08 625 struct.-based 1403.1120
canonical 802.2475
Table 4: Median times (in microseconds $\mu s$) computed over 5000 sequential repetitions to evaluate the structure-based and canonical summary statistics for different datasets.
Figure 11: Runtimes in minutes (left panel) and cumulative number of immediately rejected samples $\theta^{*}$ (right panel) of ten runs of the SMC-ABC Algorithm 1 with canonical (red lines) and structure-based summaries (black lines), respectively, as functions of the computational budget, for the setting $T_{\textrm{obs}}=200$ and $\Delta_{\textrm{obs}}=0.02$.
Appendix B Impact of the prior distribution
In this section, we investigate the proposed SBP SMC-ABC method (i.e., Algorithm 1 with structure-preserving splitting simulation, structure-based summaries and standard Gaussian proposal sampler) with different prior distributions for the model parameters. In particular, we fix $T_{\textrm{obs}}=200$ and $\Delta_{\textrm{obs}}=0.02$, and compare the estimation results based on the uniform priors (11) (cf. black lines in Figure 3 and Figure 4 of Section 4.1) with those obtained from log-normal and exponential priors, given by
$\epsilon\sim\textrm{LogN}(0,1),\quad\gamma\sim\textrm{LogN}\Bigl{(}0,\frac{1}{2}\Bigr{)},\quad\beta\sim\textrm{LogN}(0,1),\quad\sigma\sim\textrm{LogN}\Bigl{(}0,\frac{3}{4}\Bigr{)},$
$\epsilon\sim\textrm{Exp}(3),\quad\gamma\sim\textrm{Exp}\Bigl{(}\frac{1}{2}\Bigr{)},\quad\beta\sim\textrm{Exp}\Bigl{(}\frac{1}{2}\Bigr{)},\quad\sigma\sim\textrm{Exp}(1),$
respectively. Here, $\textrm{LogN}(\mu,\sigma)$ denotes a logNormal distribution where the underlying Normal has mean $\mu$ and standard deviation $\sigma$, while $\textrm{Exp}(\lambda)$ is an exponential with rate parameter $\lambda$. All investigated priors have means, medians and modes different from the true values. Moreover, they all account for the positivity of the model parameters, the log-normal and exponential ones being defined on the entire parameter space $(0,\infty)^{4}$.
Figure 12: $90$% CIs of the marginal posterior densities of the parameter vector $\theta$ (2) of the stochastic FHN model (1) obtained from Algorithm 1 with uniform (black squares), log-normal (blue circles) or exponential (red triangles) priors. The green horizontal lines and the dashed lines indicate the true parameter values and the corresponding posterior medians, respectively.
As expected, the prior distribution impacts the inferential results in the first few iterations of the algorithm, when the used computational budget is small. However, as the algorithm evolves, the impact of the chosen prior distribution on the inferential results becomes negligible. This can be observed in Figure 12, where we report the $90$% CIs of the derived ABC posterior densities for $\theta$ (2) under uniform (black squares), log-normal (blue circles) and exponential (red triangles) priors as a function of the computational budget. The CIs obtained under the different priors vary initially (see the top panels) up to a computational budget of approximately $2\cdot 10^{5}$, when they overlap. In Figure 13, we report the ABC posterior densities obtained for a computational budget of $\textrm{Nsim}_{\textrm{max}}=10^{6}$ (black solid, blue dotted and red dashed lines for uniform, log-normal and exponential priors, respectively), which hardly differ from each other, suggesting a marginal impact of the chosen prior on the inferential results.
Figure 13: Marginal posterior and prior densities of the parameter vector $\theta$ (2) of the stochastic FHN model (1) obtained from Algorithm 1 with uniform (black solid lines), log-normal (blue dotted lines) or exponential (red dashed lines) priors, when using a computational budget of $\textrm{Nsim}_{\textrm{max}}=10^{6}$. The green vertical lines indicate the true parameter values.
Appendix C Impact of the proposal kernel